OpenStudy (anonymous):

SYSTEMS Vertical Motion: If S=0.5AT^2+VoT+So, find the equation which fits the following 3 data points below . To make it easier on myself, I'm letting A=x, T=y, Vo=z, and So= w 128=0.5x(1)^2+z(1)+w 80=0.5x(2)^2+z(2)+w 0=0.5x(3)^2+z(3)+w The answer should come up to be S=-16t^2+144

OpenStudy (phi):

First step is simplify the coefficients. for example, the first equation could be 128 = 0.5x + z + w or (multiply by 2 to get rid of the decimal) 256= x + 2x + 2w

OpenStudy (anonymous):

yeah

OpenStudy (phi):

what do you get for the simplified versions of the 3 equations ?

OpenStudy (phi):

? I assume the points are (t,s) pairs. 80=0.5x(2)^2+z(2)+w means 80 = 0.5 * 2^2 * x + 2z + w there is no x^2 in there

OpenStudy (anonymous):

Ah, ok then...I got confused

OpenStudy (anonymous):

Odd, every thing I do results in a = -32, not -16 :<

OpenStudy (phi):

Yes, I get x=-32, z=0 and w=144 I would ask the teacher about this one.